# SG-LRA: Self-Generating Automatic Scoliosis Cobb Angle Measurement with Low-Rank Approximation

### v1

@@ Abstract
Automatic Cobb angle measurement from X-ray images is crucial for scoliosis screening and diagnosis. However, most existing regression-based methods and segmentation-based methods struggle with inaccurate spine representations or mask connectivity/fragmentation issues. Besides, landmark-based methods suffer from insufficient training data and annotations. To address these challenges, we propose a novel framework including Self-Generation pipeline and Low-Rank Approximation representation (SG-LRA) for automatic Cobb angle measurement. Specifically, we propose a parameterized spine contour representation based on LRA, which enables eigen-spine decomposition and spine contour reconstruction. We can directly obtain spine contour with only regressed LRA coefficients, which form a more accurate spine representation than rectangular boxes. Also, we combine LRA coefficient regression with anchor box classification to solve inaccurate predictions and mask connectivity issues. Moreover, we develop a data engine with automatic annotation and automatic selection in an iterative manner, which is trained on a private Spinal2023 dataset. With our data engine, we generate the largest scoliosis X-ray dataset named Spinal-AI2024 largely without privacy leaks. Extensive experiments on public AASCE2019, private Spinal2023, and generated Spinal-AI2024 datasets demonstrate that our method achieves state-of-the-art Cobb angle measurement performance. Our code and Spinal-AI2024 dataset are available at https://github.com/Ernestchenchen/SG-LRA and https://github.com/Ernestchenchen/Spinal-AI2024, respectively.
{IEEEkeywords}
Automatic Cobb angle measurement, self-generation pipeline, low-rank approximation representation, largest scoliosis X-ray dataset
@@ 1 Introduction
[…]
Scoliosis is an important spinal disease in human beings, especially for Scoliosis, a common spinal deformity that primarily affects adolescents, can significantly impact the health of patients [1, 2, 3]. Early detection and treatment of scoliosis can help reduce the need for surgical interventions [4]. The Cobb angle [5] is one of the most widely used clinical measures for evaluating and diagnosing the severity of scoliosis. It is defined as the angle between parallel lines of the top of the upper end vertebra (the most tilted vertebrae on the curve) and the bottom of the lower one [6]. There are three regional spinal deformities associated with different spinal segments: proximal thoracic (PT), main thoracic (MT), and thoracolumbar/lumbar (TL/L), as illustrated in Fig. 1(c).
@@
Currently, manual Cobb angle measurement on radiographic images like X-ray images by experts is still the mainstream way in medical institutions. However, it is time-consuming and often has potential errors [7, 8], since it can be affected by spinal selection and inter-observer variability, which can be observed from the distances between manual angles and angles by landmarks in Fig. 1(a). It has been validated that manual annotation of the Cobb angle exhibits fluctuations ranging from 2.8 to 8 degrees [9]. Therefore, developing a reliable automatic Cobb angle measurement method is crucial.
Traditional automatic Cobb angle measurement methods rely on hand-crafted features [10, 11], limiting model performance and generalization. In recent years, some regression-based methods use rectangular boxes as the spine representation [12, 13], which suffer from reduced accuracy when spinal blocks are wedge-shaped or irregular, as shown in Fig. 1(a). Segmentation-based methods are sensitive to image quality [14, 15], tending to predict connected or damaged spinal masks for low-quality inputs. Landmark-based methods necessitate vast amounts of training data and annotations [16, 17, 18], which are costly and time-consuming to collect and label, and are exacerbated by data scarcity due to patient privacy concerns.
@@
Parameterized Spine Representation. We propose a novel automatic Cobb angle measurement framework named SG-LRA. We combine regression-based and landmark-based techniques, by using a set of spinal landmarks to represent the bounding box of spinal segments. Inspired by low-rank approximation (LRA) parameterized contour representation in object detection [19, 20], we use LRA to learn eginespines from spine contours of training data, and reconstruct spine contours via LRA coefficients. To suppress inaccurate predictions and mask connectivity, we integrate LRA coefficient regression with anchor box classification to obtain spinal segment contours, in which sparse and dense assignments are integrated to more precisely detect spinal segments.
Privacy-Preserving Scoliosis Generation. We explore the latent diffusion model (LDM) [21] for scoliosis image synthesis. Specifically, we train LDM on our private dataset Spinal2023, utilizing a diffusion process to map raw images from pixel to latent space. During inference, the trained LDM with text-conditioned prompts can generate X-ray images. To mitigate privacy leaks [22], we employ data augmentation during LDM training, promote data balance, and introduce a privacy review [23] to exclude potentially privacy-compromising samples based on structural and pixel similarities, coupled with manual verification.
Cost-Effective Semi-Supervised Annotation. Inspired from semi-supervised pseudo-labeling [24] and multi-stage interactive annotation [25], we propose a new data engine, which drives an iterative pipeline with automatic annotation and selection. It consists of four stages: pseudo-labeling, auto-annotation, manual-assisted annotation, and privacy review. At each iteration, spine contour detection network is first re-trained using current dataset, and its trained model is used for pseudo-labeling so as to further update the dataset. This iterative pipeline culminates in our open-source dataset named Spinal-AI2024, which is the largest scoliosis X-ray dataset.
@@
  • We propose a novel low-rank approximation parameterized spine representation for automatic Cobb angle measurement, in which LRA coefficient regression and anchor box classification with sparse and dense assignments are combined to detect irregular contours of spinal segments.
  • We propose a new data engine with semi-supervised labeling and privacy review in an iterative manner, which enables the generation of large-scale scoliosis X-ray images largely without privacy leaks. To our knowledge, our open-source Spinal-AI2024 is the largest scoliosis X-ray dataset.
@@ 2.1 Scoliosis Cobb Angle Measurement
Earlier methods [10, 11, 26, 27] adopt traditional machine learning to approximate the relationship between image features and Cobb angles. However, these methods suffer from performance and generalization issues due to loss of high-resolution details through downsampling [17] or reliance on hand-crafted features. Recently, deep learning based methods [28, 29] have gained increasing attention.
Figure 2: The overall structure of our SG-LRA framework, which consists of spinal image generation, spinal contour detection, and data engine. Our private Spinal2023 dataset is used for training spinal image generation model, and also used for training initial model of spinal contour detection network. During each round of data engine, trained spinal contour detection network is adopted for pseudo-labeling, and selected data with annotations are backward used to fine-tune the spinal contour detection network. By employing the data engine to select and annotate generated images, we obtain the Spinal-AI2024 dataset.
@@
Segmentation-based methods [14, 31, 32] usually rely on U-Net [33] or its modifications [34, 15], by performing pixel-level classification and then combining pixels as spinal regions. However, these methods are sensitive to image quality, and often require pre-processing to address noise, resolution, and contrast issues [35, 32].
Landmark-based methods [36, 16, 37, 17, 38] focus on localizing spinal landmarks, which often requires vast training data. In this paper, we overcome the data scarcity issue by predicting landmarks through LRA coefficient regression, leveraging our proposed data engine with automatic annotation and selection.
@@ 2.2 Medical Image Generation
Deep learning relies heavily on numerous labeled training data [39]. In the medical image field, there is often a scarcity of annotated data due to cumbersome, time-consuming, and expensive image acquisition and annotation [40]. To alleviate this limitation, previous works use data perturbation [41] or data synthesis [42].
Generative models like generative adversarial networks (GANs) [43, 44] have demonstrated the capability to produce realistic, high-resolution images [44, 45, 46, 47]. Recent studies [22] indicate that diffusion models [21] possess higher generation capacity than GANs. Packhäuser et al. [23] utilized LDM to generate chest X-ray images, and proposed a sampling strategy that safeguards the privacy of sensitive biometric information. However, sharing sensitive data under stringent privacy regulations remains a critical challenge in medical research [48, 49]. In this paper, we also leverage LDM to generate scoliosis X-ray images. Additionally, we propose a privacy-auditing module combined with pseudo-label training, utilizing a comprehensive privacy similarity mechanism to filter out images that may expose patient privacy.
@@ 2.3 Semi-Supervised Labeling
Semi-supervised labeling methods can be categorized into consistency regularization and pseudo-labeling approaches. Consistency regularization methods [50, 51] rely on modality-specific augmentation for regularization. Pseudo-labeling methods [52] predict labels for unlabeled data and train the model in a supervised manner.
@@ 3 Methodology
Also, we use Spinal2023 to train an initial model of spinal contour detection network. This network adopts ResNet50 [57] as backbone, and uses feature pyramid network (FPN) [58] to extract multi-scale features. Then, it uses regression and classification branches to predict LRA coefficients and positive anchor boxes, respectively. In both branches, sparse and dense assignments are combined to improve the detection performance. Finally, we select LRA coefficients corresponding to positive anchor boxes to reconstruct spinal segment contours.
@@
3.1 Spinal Contour Detection via LRA
@@ 3.1.1 Parameterized Spine Representation
where $\tilde{\mathbf{a}}_{i}=\mathbf{U}_{M}\mathbf{c}_{i}=[\mathbf{u}_{1},\mathbf{u}_{2},\cdots,\mathbf{u}_{M}]\mathbf{c}_{i}$ is the approximation of $\mathbf{a}_{i}$. We refer to $\mathbf{u}_{1},\ldots,\mathbf{u}_{M}$ as eigen-spines, since they are eigen-vectors of the matrix $\mathbf{AA^{T}}$ and can be viewed as principal spinal segment shapes.
@@
where $\mathbf{c}\in\mathbb{R}^{M\times 1}$ is the LRA coefficient vector.
@@ 3.1.2 Spinal Contour Detection Network
where $l_{1}$ denotes smooth-L1 loss, $ps$ denotes the positive sample region obtained by incorporating sparse and dense assignments, and $\mathds{1}$ denotes indicator function, outputting 1 when the condition is valid (the $i$-th point is within $ps$ in Eq. (7)) and outputting 0 otherwise. After combining the outputs of regression and classification branches, we can obtain a LRA coefficient matrix $\mathbf{C}_{K}\in\mathbb{R}^{M\times K}$. The predicted spinal segment locations are calculated as
@@
Algorithm 1 The overall procedure of our data engine.
@@
1:  Train spinal contour detection network on $D_{spinal2023}$ and obtain initial parameters $\Theta^{0}$. Set $i\leftarrow 1$.
2:  while $V_{selected}$ is not converged do
3:     Pseudo-label samples in $D_{selected}$ using $\Theta^{i-1}$ with $\tau_{c}$-based positive spinal segment instance selection.
4:     Update $V_{selected}$ by auto-annotation, manual-assisted annotation, and privacy review.
5:     $D_{SpinalAI2024}\leftarrow\{s^{j}|s^{j}\in D_{selected},v^{j}=1\}$.
6:     Fine-tune network from $\Theta^{i-1}$ using samples from $D_{spinal2023}\cup D_{SpinalAI2024}$, and obtain new parameters $\Theta^{i}$. Set $i\leftarrow i+1$.
7:  end while
8:  Return $D_{SpinalAI2024}$.
3.2 Data Engine with Dataset Generation
Our data engine is composed of iteratively four stages: pseudo-labeling, auto-annotation, manual-assisted annotation, and privacy review. Before starting the data engine, we use Stable Diffusion [21] to generate an initial set of spinal X-ray images, as shown in Algorithm 1.
@@ 3.2.1 Pseudo-Labeling
At the $i$-th iteration, to enable pseudo-labeling, we first train our spinal contour detection network on currently available labeled data composed of $D_{spinal2023}$ and $D_{SpinalAI2024}$, in which the trained model parameters are denoted as $\Theta^{i-1}$. Then, we use $\Theta^{i-1}$ to label samples in $D_{selected}$. For the $j$-th sample $s^{j}$, the network predicts the locations of $K$ spinal segments $[\mathbf{a}_{1}^{j},\mathbf{a}_{2}^{j},\cdots,\mathbf{a}_{K}^{j}]$. To filter inaccurate predictions, we use a confidence threshold $\tau_{c}$ to select positive spinal segment instances for $s^{j}$:
@@
Table 1: The statistical characteristics of different datasets.
train test train test train test
Number of Images 481 128 700 204 16000 4000
Height of Images (Pixel) [973-3755] [1159-2880] [184-833] [227-802] [512-512] [512-512]
Width of Images (Pixel) [355-1427] [427-1386] [135-457] [147-415] [512-512] [512-512]
Cobb Angle (°) [5.96-90.0] [8.21-80.99] [3.22-74.26] [5.16-67.18] [0.0-90.0] [0.0-90.0]
Scoliosis Severity Num Avg Num Avg Num Avg Num Avg Num Avg Num Avg
Not ($<$ 10°) 17 8.35 1 8.21 91 7.82 20 8.24 1123 8.27 278 7.86
Severe ($>$ 45°) 152 60.89 43 56.98 87 55.87 23 56.1 1101 60.41 359 59.05
@@ 3.2.4 Privacy Review
After the above selection, we can obtain currently labeled generated dataset $D_{SpinalAI2024}$ using the selection index set $V_{selected}$. We then merge $D_{SpinalAI2024}$ with the existing $D_{spinal2023}$ to fine-tune the spinal contour detection network. We continue to generate pseudo-labels at the next iteration, and repeat this process until $V_{selected}$ converges.

### v2

# CurvNet: Latent Contour Representation and Iterative Data Engine for Curvature Angle Estimation

@@ Abstract
Curvature angle is a quantitative measurement of a curve, in which Cobb angle is customized for spinal curvature. Automatic Cobb angle measurement from X-ray images is crucial for scoliosis screening and diagnosis. However, most existing regression-based and segmentation-based methods struggle with inaccurate spine representations or mask connectivity and fragmentation issues. Besides, landmark-based methods suffer from insufficient training data and annotations. To address these challenges, we propose a novel curvature angle estimation framework named CurvNet including latent contour representation based contour detection and iterative data engine based image self-generation. Specifically, we propose a parameterized spine contour representation in latent space, which enables eigen-spine decomposition and spine contour reconstruction. Latent contour coefficient regression is combined with anchor box classification to solve inaccurate predictions and mask connectivity issues. Moreover, we develop a data engine with image self-generation, automatic annotation, and automatic selection in an iterative manner. By our data engine, we generate a clean dataset named Spinal-AI2024 without privacy leaks, which is the largest released scoliosis X-ray dataset to our knowledge. Extensive experiments on public AASCE2019, our private Spinal2023, and our generated Spinal-AI2024 datasets demonstrate that our method achieves state-of-the-art Cobb angle estimation performance. Our code and Spinal-AI2024 dataset are available at https://github.com/Ernestchenchen/CurvNet and https://github.com/Ernestchenchen/Spinal-AI2024, respectively.
keywords:
Object detection , Image generation , Curvature angle estimation , Latent contour representation , Iterative data engine , Largest released scoliosis X-ray dataset
journal: Pattern Recognition
@@ 1 Introduction
Automatic measurement of curvature angle is important for curved objects. Due to critical role of spine in human body, Cobb angle estimation of scoliosis, a common spinal deformity disease in human beings, has gained increasing attention in the field of artificial intelligence and computer vision [1]. Early detection and treatment of scoliosis can help reduce the need for surgical interventions [2]. The Cobb angle is one of the most widely used clinical measures for evaluating and diagnosing the severity of scoliosis. It is defined as the angle between parallel lines of the top of the upper end vertebra (the most tilted vertebrae on the curve) and the bottom of the lower one [3]. There are three regional spinal deformities associated with different spinal segments: proximal thoracic (PT), main thoracic (MT), and thoracolumbar/lumbar (TL/L), as illustrated in Fig. 1(c).
Currently, manual Cobb angle measurement on radiographic images like X-ray images by experts is still the mainstream way in medical institutions. However, it is time-consuming and often has potential errors, since it can be affected by spinal selection and inter-observer variability, which can be observed from the distances between manual angles and angles by landmarks in Fig. 1(a). It has been validated that manual annotation of the Cobb angle exhibits fluctuations ranging from 2.8 to 8 degrees [4]. Therefore, developing a reliable automatic Cobb angle measurement method is crucial.
Traditional automatic Cobb angle measurement methods rely on hand-crafted features [5], limiting model performance and generalization. In recent years, some regression-based methods use rectangular boxes as the spine representation [6], which suffer from reduced accuracy when spinal blocks are wedge-shaped or irregular, as shown in Fig. 1(a). Segmentation-based methods are sensitive to image quality [7, 8], tending to predict connected or damaged spinal masks for low-quality inputs. Landmark-based methods necessitate vast amounts of training data and annotations [9, 10], which are costly and time-consuming to collect and label, and are exacerbated by data scarcity due to patient privacy concerns.
@@
Parameterized Spine Representation. We propose a novel automatic Cobb angle measurement framework named CurvNet. We combine regression-based and landmark-based techniques, by using a set of spinal landmarks to represent the bounding box of spinal segments. Inspired by low-rank approximation parameterized contour representation in object detection [11], we use latent space to learn eigen-spines from spine contours of training data, and reconstruct spine contours via latent contour coefficients. To suppress inaccurate predictions and mask connectivity, we integrate latent contour coefficient regression with anchor box classification to obtain spinal segment contours, in which sparse and dense assignments are integrated to more precisely detect spinal segments.
Privacy-Preserving Scoliosis Generation. We explore the latent diffusion model (LDM) [12] for scoliosis image self-generation. Specifically, we train LDM on our private dataset Spinal2023, utilizing a diffusion process to map raw images from pixel to latent space. During inference, the trained LDM with text-conditioned prompts can generate X-ray images. To mitigate privacy leaks [13], we employ data augmentation during LDM training, promote data balance, and introduce a privacy review [14] to exclude potentially privacy-compromising samples based on structural and pixel similarities, coupled with manual verification.
Cost-Effective Semi-Supervised Annotation. Inspired from semi-supervised pseudo-labeling [15] and multi-stage interactive annotation [16], we propose a new data engine, which drives an iterative pipeline with automatic annotation and selection. It consists of four stages: pseudo-labeling, auto-annotation, manual-assisted annotation, and privacy review. At each iteration, spine contour detection network is first re-trained using current dataset, and its trained model is used for pseudo-labeling so as to further update the dataset. This iterative pipeline culminates in our open-source clean dataset named Spinal-AI2024, which is the largest released scoliosis X-ray dataset to our knowledge.
@@
  • 1.
    We propose a novel low-rank approximation parameterized spine representation for curvature angle estimation, in which latent contour coefficient regression and anchor box classification with sparse and dense assignments are combined to detect irregular contours of spinal segments.
  • 2.
    We propose a new data engine with semi-supervised labeling and privacy review in an iterative manner, which enables the generation of large-scale clean scoliosis X-ray images without privacy leaks. To our knowledge, our open-source Spinal-AI2024 is the largest released scoliosis X-ray dataset.
  • 3.
@@ 2.1 Scoliosis Cobb Angle Measurement
Earlier methods [17, 5] adopt traditional machine learning to approximate the relationship between image features and Cobb angles. However, these methods suffer from performance and generalization issues due to loss of high-resolution details through downsampling [10] or reliance on hand-crafted features. Recently, deep learning based methods [18] have gained increasing attention.
@@
Segmentation-based methods [7, 22] usually rely on U-Net [23] or its modifications [24, 8], by performing pixel-level classification and then combining pixels as spinal regions. However, these methods are sensitive to image quality, and often require pre-processing to address noise, resolution, and contrast issues [22].
Landmark-based methods [25, 9, 26, 10, 27] focus on localizing spinal landmarks, which often requires vast training data. In this paper, we overcome the data scarcity issue by predicting landmarks through latent contour coefficient regression, leveraging our proposed data engine with automatic annotation and selection.
@@ 2.2 Medical Image Generation
Deep learning relies heavily on numerous labeled training data. In the medical image field, there is often a scarcity of annotated data due to cumbersome, time-consuming, and expensive image acquisition and annotation. To alleviate this limitation, previous works use data perturbation or data synthesis.
Generative models like generative adversarial networks (GANs) [28] have demonstrated the capability to produce realistic, high-resolution images. Recent studies [13] indicate that diffusion models [12] possess higher generation capacity than GANs. Packhäuser et al. [14] utilized latent diffusion model (LDM) to generate chest X-ray images, and proposed a sampling strategy that safeguards the privacy of sensitive biometric information. However, sharing sensitive data under stringent privacy regulations remains a critical challenge in medical research [29]. In this paper, we also leverage LDM to generate scoliosis X-ray images. Additionally, we propose a privacy-auditing module combined with pseudo-label training, utilizing a comprehensive privacy similarity mechanism to filter out images that may expose patient privacy.
@@ 2.3 Semi-Supervised Labeling
Semi-supervised labeling methods can be categorized into consistency regularization and pseudo-labeling approaches. Consistency regularization methods [30] rely on modality-specific augmentation for regularization. Pseudo-labeling methods [31] predict labels for unlabeled data and train the model in a supervised manner.
@@
Figure 2: The overall structure of our CurvNet framework, which consists of spinal image generation, spinal contour detection, and data engine. Our private Spinal2023 dataset is used for training spinal image generation model, and also used for training initial model of spinal contour detection network. During each round of data engine, trained spinal contour detection network is adopted for pseudo-labeling, and selected data with annotations are backward used to fine-tune the spinal contour detection network. By employing the data engine to select and annotate generated images, we obtain the Spinal-AI2024 dataset.
@@ 3 Methodology
Also, we use Spinal2023 to train an initial model of spinal contour detection network. This network adopts ResNet50 [36] as backbone, and uses feature pyramid network (FPN) [37] to extract multi-scale features. Then, it uses regression and classification branches to predict latent contour representationand positive anchor boxes, respectively. In both branches, sparse and dense assignments are combined to improve the detection performance. Finally, we select latent contour coefficientscorresponding to positive anchor boxes to reconstruct spinal segment contours.
@@
3.1 Spinal Contour Detection via Latent Contour Representation
@@ 3.1.1 Parameterized Spine Representation
where $\tilde{\mathbf{a}}_{i}=\mathbf{U}_{M}\mathbf{c}_{i}=[\mathbf{u}_{1},\mathbf{u}_{2},\cdots,\mathbf{u}_{M}]\mathbf{c}_{i}$ is the approximation of $\mathbf{a}_{i}$.
We refer to $\mathbf{u}_{1},\ldots,\mathbf{u}_{M}$ as eigen-spines, since they are eigen-vectors of the matrix $\mathbf{AA^{T}}$ and can be viewed as principal spinal segment shapes.
@@
where the eigen-spine matrix $\mathbf{U}_{M}\in\mathbb{R}^{2N\times M}$ projects $\mathbf{a}\in\mathbb{R}^{2N\times 1}$ into the latent space, and $\mathbf{c}\in\mathbb{R}^{M\times 1}$ is the latent contour coefficient vector.
@@ 3.1.2 Spinal Contour Detection Network
where $l_{1}$ denotes smooth-L1 loss, $ps$ denotes the positive sample region obtained by incorporating sparse and dense assignments, and $\mathds{1}$ denotes indicator function, outputting 1 when the condition is valid (the $i$-th point is within $ps$ in Eq. (7)) and outputting 0 otherwise. After combining the outputs of regression and classification branches, we can obtain a latent contour coefficient matrix $\mathbf{C}_{K}\in\mathbb{R}^{M\times K}$. The predicted spinal segment locations are calculated as
@@
Algorithm 1 The overall procedure of our iterative data engine.
@@
1: Train spinal contour detection network on $D_{spinal2023}$, and obtain initial parameters $\Theta^{0}$.
2: Set $i\leftarrow 1$.
3:while $V_{selected}$ is not converged do
4:  Pseudo-label samples in $D_{selected}$ using $\Theta^{i-1}$ with $\tau_{c}$-based positive spinal segment instance selection.
5:  Update $V_{selected}$ by auto-annotation, manual-assisted annotation, and privacy review.
6:  $D_{spinalAI2024}\leftarrow\{s^{j}|s^{j}\in D_{selected},v^{j}=1\}$.
7:  Fine-tune network from $\Theta^{i-1}$ using samples from $D_{spinal2023}\cup D_{spinalAI2024}$, and obtain new parameters $\Theta^{i}$.
8:  Set $i\leftarrow i+1$.
9:end while
10: Return $D_{spinalAI2024}$.
3.2 Iterative Data Engine with Image Self-Generation
Our data engine is composed of iteratively four stages: pseudo-labeling, auto-annotation, manual-assisted annotation, and privacy review. Before starting the data engine, we use Stable Diffusion [12] to generate an initial set of spinal X-ray images, as shown in Algorithm 1. We use an early stopping strategy for the iterative process from Step 3 to Step 9, in which the iteration stops when the number of newly selected samples changes by less than 1% over two consecutive rounds, or after six iterations.
@@ 3.2.1 Pseudo-Labeling
At the $i$-th iteration, to enable pseudo-labeling, we first train our spinal contour detection network on currently available labeled data composed of $D_{spinal2023}$ and $D_{spinalAI2024}$, in which the trained model parameters are denoted as $\Theta^{i-1}$. Then, we use $\Theta^{i-1}$ to label samples in $D_{selected}$. For the $j$-th sample $s^{j}$, the network predicts the locations of $K$ spinal segments $\mathbf{a}_{1}^{j},\mathbf{a}_{2}^{j},\cdots,\mathbf{a}_{K}^{j}]$. To filter inaccurate predictions, we use a confidence threshold $\tau_{c}$ to select positive spinal segment instances for $s^{j}$:
@@ 3.2.4 Privacy Review
After the above selection, we can obtain currently labeled generated clean dataset $D_{spinalAI2024}$ using the selection index set $V_{selected}$. We then merge $D_{spinalAI2024}$ with the existing $D_{spinal2023}$ to fine-tune the spinal contour detection network. We continue to generate pseudo-labels at the next iteration, and repeat this process until $V_{selected}$ converges.
Table 1: The statistical characteristics of different datasets.
train test train test train test
Number of Images 481 128 700 204 16000 4000
Height of Images (Pixel) 973-3755 1159-2880 184-833 227-802 512-512 512-512
Width of Images (Pixel) 355-1427 427-1386 135-457 147-415 512-512 512-512
Cobb Angle (°) 5.96-90.0 8.21-80.99 3.22-74.26 5.16-67.18 0.0-90.0 0.0-90.0
Scoliosis Severity Num Avg Num Avg Num Avg Num Avg Num Avg Num Avg
Not (<< 10°) 17 8.35 1 8.21 91 7.82 20 8.24 1123 8.27 278 7.86
Severe (>> 45°) 152 60.89 43 56.98 87 55.87 23 56.1 1101 60.41 359 59.05